Project

General

Profile

NDOS Commissioning Meeting 2011/05/23

Attendance: Jon, Gary, Nick, Zeb, Tom, Jarek, Sue, Rick, Greg, Peter, Gavin, Minerba, Mat, Kurt, Leon

Agenda:

  1. NDOS Status and Channel Recovery Plans
  2. DAQ/Electronics Update & Plans
  3. DCS Update & Plans
  4. AOB (All)

NDOS Status and Channel Recovery Plans

  • The production dB will be shutdown tomorrow morning at 8 am in order for the machine to be moved. This means the hardware database and online logbook will be down during this time. Kurt and Jon will switch the DAQ to use the dev. dB at 3 pm today.
  • Mat is still digesting the cable-ectomy data. Mat has made plots comparing post cable-ectomy DSO scans to pre cable-ectomy DSO scan, which confirms what was shown at the last meeting. He'd like to do another ~15 cable-ectomies (diff. FEBs). That will happen either Tues. or Wed.
  • Humidity issues:
    • Mat: any contact with Bill Moorehouse. Jon: spoke with Bill and he was going to look into either purchasing another (residential?) dehumidifier unit to put into the building, or buy an industrial unit. He needed to speak with Rick T. about this. Rick then followed up with an email to Jon stating that he is working on getting new door thresholds installed at NDSB. The plan is to see if this helps first before any new dehumidifiers.
    • UMN: use DataCheck to correlate humidity to noise? Will work on getting something ready for Thursday's meeting.
    • Gary: do we have a spec for NDSB? It would be interesting to see if we're ok for the FD? Jon will follow up with Steve Dixon.

DAQ/Electronics Update & Plans

  • Denis is working this morning on getting pedestal scanner to work with latest change to DCM names.
  • We've been without BNEVBs 1-4 since Friday. Andrew tried rebooting them on Friday; found them physically turned off. They were turned back on, but soon after there were problems so they were left out of the DAQ system. Still investigating.
  • Weekend went ok; a few hiccups but shifters recovered.
  • Neil and Rick are working on the timing link calibration firmware that will remove delays in links. Found and fixed a few bugs.
  • Last week Rick and Ron worked on SERDES timing link unlock error.
  • Firmware will come out hopefully in a few days.

DCS Update & Plans

  • Leon: no news. Once new firmware is in with unlock error bug fix, we can try reading out detector.
  • Sue: still need to resolve conflict at startup and how to timeshare with timing system.

AOB (All)